Train A leaves the station traveling east at 45 mph. Three hours later train B leaves the same station traveling west at 55 mph. How long after train B leaves the station until the trains are 610 miles apart?
Answer by Alan3354(21605)

Train A leaves the station traveling east at 45 mph. Three hours later train B leaves the same station traveling west at 55 mph. How long after train B leaves the station until the trains are 610 miles apart?
===========================
Going opposite directions, their speeds relative to each other is the sum, 45+55 mph, = 100 mph.
Train is 45*3 miles away = 135 miles.
610 - 135 = 475 miles
475/100 = 4.75 hours
= 4 hours 45 minutes